# Personalizing Agent Privacy Decisions via Logical Entailment

### v2

@@ 7.1. Personalization and Privacy Norms
Our results in Section 6 suggest that effective personalization of LLM agents requires grounding judgments in prior user privacy judgments, as general norms are insufficient. In this paper, we consider user judgments to prior requests as proxies for the user’s privacy preferences. However, this finding does not invalidate use general norms in LLMs. Aligning models with relevant contextual norms remains crucial for many data-sharing scenarios (mireshghallah2024can; shao2024privacylens). Our work highlights that when personalization is a crucial component for making privacy decisions, then the LLM’s privacy judgments should be grounded in user-specific preferences rather than general privacy norms.
7.2. Personalization Challenges
Personalizing LLM-based agents introduces several challenges that must be considered.
Contradictory privacy preferences. The first challenge, highlighted by the error analysis in Section 6.3, arises from contradictions within the user’s own history of prior judgments. For example, the user says it is inappropriate to share their banking information with a third party if their data was confidential, but deems it appropriate if there is no safeguard on the data. In such scenarios, the agent may execute the entailment logic correctly but still reach an incorrect privacy judgment. Consequently, the agent must (1) detect these user privacy judgment contradictions and (2) resolve them, typically by escalating the request to the user. The fundamental difficulty lies in distinguishing whether an incorrect privacy judgment stems from an error in the entailment process or an inconsistency in the user’s prior privacy judgments.
Evolving privacy preferences. The second challenge involves the dynamic nature of privacy preferences as they can change overtime. For example, a user can become more conservative about sharing their data as they get older. Hence, a system similar to ARIEL requires a mechanism to detect when the user’s privacy preferences are shifting and determine which preferences must be updated. One potential mechanism could trade-off exploration and exploitation by escalating incoming requests to the user every once in a while to confirm the agent’s decision.
7.3. User Agency in Personal AI Agents
While personalized agents reduce the burden of privacy management, they must not remove user agency (mitchell2025fully). This observation presents a tradeoff where the agent automates clear decisions but escalates cases where the judgment cannot be easily inferred by the user’s prior judgments. ARIEL implements this via an ’undetermined’ option, and our results confirm that identifying these non-inferable cases drives ARIEL’s judgment accuracy gains. Furthermore, by providing transparent reasoning traces for user auditing, ARIEL successfully balances automation with necessary human oversight. Future work could operationalize this transparency through a dedicated user interface, enabling users to visualize the agent’s logic and intervene in decision-making when necessary.
@@ 7.4. Limitations
There are three main limitations involving our evaluation in this work. The first is that our evaluation datasets simulate prior requests by partitioning the users requests into two pairwise disjoint sets, where one set stores already observed requests while the other contains new requests. While we believe this reasonably approximates our problem setup, it does not incorporate user decision-making over time. Unfortunately, to our knowledge, there are no existing publicly-available datasets that incorporates user privacy preferences over time. An interesting future work could elicit user privacy preferences over a period of time, rather than giving a one-time questionnaire to participants. The second is our evaluation does not consider cold-start cases where the user has not made prior judgments on requests– essentially, no privacy preferences. While falling back to general privacy norms is a potential solution, our results indicate they are often insufficient for personalized privacy decisions. An alternative solution could be to develop minimal question sets to elicit the user’s privacy preferences (guo2025not; tran2025understanding; zhang2025towards). However, as our work focuses on inferring privacy judgments on behalf of users, we leave this elicitation as a future work. Lastly, our work assumes a non-adversarial setting where the incoming requests are not maliciously designed. While this limits applicability in open environments, it is appropriate for controlled ecosystems where input sources are authenticated and vetted.
@@ 8. Conclusion
In this work, we explored personalization of LLM-based agents for making privacy judgments. We identified that general privacy norms are insufficient for effective personalization, and LLM-based reasoning with ICL does not guarantee judgments faithfully grounded on user’s prior privacy judgments. We addressed these limitations by designing ARIEL to ensure judgments are grounded in user’s prior judgments with clear and traceable reasoning that the user can reference. Although ARIEL is intended to accurately make privacy judgments on behalf of users, it is designed with user agency as a key component of the framework. Our evaluations demonstrate that ARIEL can improve appropriate judgments by reducing the F1 score error of appropriate judgments by up to 39.1% over LLM reasoning. Overall, our findings suggest that jointly leveraging LLMs with logical entailment is a promising direction for designing agentic frameworks for personalized privacy judgments.
@@ Acknowledgements.
We would like to thank Adrià Gascón, Sarah Meiklejohn, Eugene Bagdasarian, Daniel Ramage, Zhichen Zhao, and Ku Kogan for their constructive discussions and support.

### v3

@@ 7.3. Personalization and Privacy Norms
Our results in Section 6 suggest that effective personalization of LLM agents requires grounding judgments in prior user privacy judgments, as general norms are insufficient. However, this finding does not invalidate alignment of LLMs with general contextual norms, as this alignment remains crucial for many data-sharing scenarios (mireshghallah2024can; shao2024privacylens). Our work highlights that when personalization is a crucial component for making privacy decisions, then the LLM’s privacy judgments should be grounded in user-specific preferences rather than general privacy norms.
@@ 7.4. Limitations
We identify four main limitations involving our work.
Ontology Mapping and Generation
ARIEL generates generalizable ontologies for individual fields (e.g., data type) based on the user’s prior privacy judgments. However, in practice, the ontology of one field may depend on the values of others. For instance, data sensitivity can vary based on the recipient: a user may consider their income more sensitive than SSN when sharing with a parent, but the sensitivity is reversed when sharing with a friend. Our results in Section 6.1 demonstrate that even without explicitly modeling these conditional dependencies, ARIEL accurately infers privacy judgments. We leave the incorporation of context-conditional hierarchies for future work.
Evaluation Datasets
Our evaluation datasets simulate prior requests by partitioning the user’s judgments on requests into two pairwise disjoint sets, where one set stores already observed requests while the other contains new requests. While we believe this reasonably approximates our problem setup, it does not incorporate user decision-making over time. Unfortunately, to our knowledge, there are no existing publicly-available datasets that incorporate user privacy preferences over time. An interesting direction for future work could elicit user privacy preferences over a period of time, rather than giving a one-time questionnaire to participants.
Although we believe our evaluation datasets accurately reflect our problem setup, they were not originally framed as AI agents acting on the users’ behalf. This distinction is notable, as prior work indicates that privacy preferences can shift when information-sharing decisions are delegated to a personal AI agent (guo2025not). While recent studies have begun eliciting user privacy preferences regarding AI agents and chatbots (guo2025not; tran2025understanding; zhang2025autonomy; liu2025prevalence; wu2025towards; zufferey2025ai), these datasets are either (1) not publicly available, or (2) restricted to narrowly scoped scenarios. Consequently, this emerging field would benefit significantly from future work that appropriately releases open datasets on user privacy preferences in agentic data-sharing contexts.
Cold-start Cases
Our evaluation does not consider cold-start cases where the user has not made prior judgments on requests– essentially, no prior privacy preferences. Handling these cold-start cases is an important consideration when designing personalized agents. While falling back to general privacy norms is a potential solution, our results demonstrate they are often insufficient for personalized privacy decisions. An alternative solution could be to develop the minimal set of questions to elicit the user’s privacy preferences (guo2025privi) while offsetting user burden. We leave evaluation on cold-start cases as future work.
Non-adversarial Setting
Lastly, our work assumes a non-adversarial setting which could limit the applicability of ARIEL in open environments. However, ARIEL is designed as a specialized module focused on agent privacy decisions, functioning within a larger system that authenticates inputs. An example is Google Play Store where app developers are verified and users provide feedback on problematic apps, which prevents requests to the agent from being maliciously designed. Thus, defense against adversarial attacks, such as prompt injection (greshake2023not), is delegated to separate system components that focus specifically on these attacks (meng2025cellmate; tsai2025contextual). Additionally, ARIEL operates on limited structured input containing just five fields, inherently reducing the risk of adversarial attacks.
@@ 8. Conclusion
In this work, we explored personalization of LLM-based agents for making privacy judgments. We identified that general privacy norms are insufficient for effective personalization, and LLM-based reasoning with ICL does not guarantee that judgments are faithfully grounded in the user’s prior privacy judgments. We addressed these limitations by designing ARIEL to ensure judgments are grounded in the user’s prior judgments with clear and traceable reasoning that the user can reference. Our evaluations demonstrate that ARIEL can reduce the F1 score error for appropriate judgments by up to 40.6% over LLM reasoning. Overall, our findings suggest that jointly leveraging LLMs with logical entailment is a promising direction for designing agentic frameworks for personalized privacy judgments.
@@ Acknowledgements.
We would like to thank Adrià Gascón, Sarah Meiklejohn, Eugene Bagdasarian, Daniel Ramage, Zhichen Zhao, and Ku Kogan for their constructive discussions and support. The authors used generative AI-based tools to revise the text, improve flow and correct any typos, grammatical errors, and awkward phrasing. This work is supported by NSF award number DGE-1842487 (J.F.) and the REAL@USC-Meta center (J.F., M.A.).